OpenStudy (anonymous):

Of the four expressions presented here, which two are equivalent? Expression I: 6n + 5 Expression II: 2n + 5 + n + 3n Expression III: 5n + 5 Expression IV: 11n I and II II and III III and IV I and IV

OpenStudy (anonymous):

Its Either B Or D. My Choice Was D But I Wanna Make Sure

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@pooja195

OpenStudy (welshfella):

simplify the second one. bring like terms together 2n + 3n + n = 5 = ?

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

welshguy. so it is B Not D?

OpenStudy (welshfella):

what is 2n + 3n + n

OpenStudy (anonymous):

I DUNNO THATS WHY IM ASKING XD

OpenStudy (anonymous):

5n + n?

OpenStudy (welshfella):

you are getting there what is 5n + n ?

OpenStudy (anonymous):

5n.

OpenStudy (welshfella):

if i have 5 cats and i get another cat how many acts have i got?

OpenStudy (welshfella):

* cats lol

OpenStudy (anonymous):

2n + 5 + n + 3n is what it is.. not what ur putting

OpenStudy (anonymous):

so if i simplifiy what i put it becomes 5n + 5 + n

OpenStudy (welshfella):

suppose n stands for a cat 5 cats + 1 cat = 6 cats 5n + n = 6n

OpenStudy (welshfella):

you can add all the n's together.

OpenStudy (anonymous):

but it doesnt tell you n = 1.. you cant add n and assume its gonna make 5n 6n

OpenStudy (welshfella):

n is the same as 1n

OpenStudy (welshfella):

n + n = 2n n + 5n = 6n

OpenStudy (welshfella):

1n + 5n = 6n

OpenStudy (anonymous):

so then its A?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

how u said its 6n alone not 6n + 5n

OpenStudy (anonymous):

Expression I: 6n + 5

OpenStudy (anonymous):

A = expression 1 and expression 2.

OpenStudy (welshfella):

you can only add LIKE terms n and 2n and 3n all have n in them they are like terms 5 is different

OpenStudy (welshfella):

yes thats right

OpenStudy (anonymous):

yeah. 5 is not a like term. so where did the 5 come from.

OpenStudy (welshfella):

It didn't come from anywhere its just there

OpenStudy (anonymous):

how is it there. how did it get there.

OpenStudy (anonymous):

you said it was 6n not 6n + 5

OpenStudy (anonymous):

so how did the 5 get there...

OpenStudy (welshfella):

It might help to think of n as a box of apples and the 5 as 5 apples so 2n + 5 is